OpenStudy (anonymous):

Describe the nature of the roots for this equation x^2-2x+1=0 and this one x^2+3x-7=0

OpenStudy (anonymous):

Hm, do you know the quadratic formula?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

well let's begin with the first one. \[ x^2 -2x + 1 = 0 \\ (x - 1)(x-1) = 0 \\ (x-1)^2 = 0 \;// \sqrt{}\\ x-1 = 0 \\ x = 1 \]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so would it be one real double root?

OpenStudy (anonymous):

For the second I don't think you can do this without the formula. Because 7 is a prime number, we can't simplify it that way. The formula looks like that \[ ax^2 + bx + c = 0 \\ x_{1,2} = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a } \] I can write down a proof if you wish, it's not complicated. And yes, there is only one x value that makes it be equal to 0.

OpenStudy (anonymous):

Perhaps a better explanation to what I meant up there. Since the only 2 numbers that can multply up to 7 are 1 and 7 and writing this \[ (x-1)(x+7) = x^2 -x +7x - 7 = x^2 +6x - 7\\ (x+1)(x-7) = x^2 + x +-7x - 7 = x^2 -6x - 7 \] we can only construct those two...